OpenStudy (anonymous):

[4.07] Choose the slope-intercept equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6. y = 1 over 3x - 8 y = -3x + 12 y = 3x - 24 y = -1 over 3x - 4

OpenStudy (hoblos):

product of slopes of 2 perpendicular lines = -1 let m be the slope of the line then 3 x m = -1 m= -1/3 now use the formula y+6 = (-1/3)(x-6) y+6 = (-1/3)x +2 y = (-1/3)x -4